Circular No. 9100

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1722 AQUILAE = NOVA AQUILAE 2009 K. Nishiyama (Kurume, Fukuoka-ken, Japan) and F. Kabashima (Mi-

yaki-cho, Saga-ken, Japan) report their discovery of a possible nova (mag 10.9) on two 40-s unfiltered CCD frames (limiting mag 13.3) taken on Dec. 14.40 UT using a 105-mm f/4 camera lens. Confirming frames (limiting mag 18.4) taken around Dec. 14.408 with a 0.40-m reflector yield mag 10.9 and the following position for the variable: α = 19^{h}14^{m}09^{s}.73, δ = +15^{o}16′34′′.7 (equinox 2000.0). Following posting on the CBAT’s unconfirmed-objects webpage, P. Corelli (Pagnacco, Italy, 0.45-m reflector) writes that an unfiltered CCD exposure from Dec. 14.79 yields position end figures 09^{s}.69, 34′′.0; Corelli adds that nothing is visible at this position on a Digitized Sky Survey plate (limiting red mag ≈ 21.5). Additional magnitudes for the new object: Dec. 6.417, [13.7 (Nishiyama and Kabashima); 7.412, [13.6 (Nishiyama and Kabashima); 14.710, 11.4 (W. Vollmann, Vienna, Austria; visual); 14.791, 10.8 (Corelli); 15.700, 10.4 (K. Hornoch, Ondrejov, Czech Republic; visual). Spectroscopic confirmation of this variable as a nova was reported by K. Kinugasa, H. Takahashi, S. Honda, H. Taguchi, and O. Hashimoto with the Gunma Astronomical Observatory 1.5-m telescope (+ GLOWS; range 420– 800 nm; resolution 1.0 nm) on Dec. 15.4 UT; and by U. Munari, Istituto Nazionale di Astrofisica, Padova Astronomical Observatory; and P. Valisa, Asiago Novae and Symbiotic Stars (ANS) collaboration, with the Varese 0.6-m telescope (range 390–860 nm) on Dec. 15.8. Munari and Valisa find the nova to be of the Fe II type, with Balmer and Fe II lines in emission and displaying P-Cyg profiles, adding that strong interstellar lines and diffuse interstellar bands suggest a high reddening. Both groups find the FWHM of the Hα emission to be ∼ 1000–1025 km/s; the Balmer series and Fe II (multiplets 42, 49) show P-Cyg profiles. Munari notes that the continuum is characterized by the deep absorption lines characteristic of novae close to their maximum brightness. The following magnitudes were measured by Munari and his ANS-collaboration colleagues G. L. Righetti, S. Dallaporta, and G. Cherini (using Tycho-2 field stars around the nova, transformed to Johnson magnitudes following Bessell 2000, PASP 112, 961): Dec. 15.74, V = 10.43, B–V = +1.43; 16.70, V = 10.32, B–V = +1.59; 17.69, V = 10.29, B–V = +1.70; 18.69, V = 10.67, B–V = +1.46.

Additional details are given on CBETs 2075 and 2076. E. Kazarovets reports that the GCVS team has assigned the designation V1722 Aql to this nova.

2009 December 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green